Due Tomorrow at 11 PM CDT Aa Aa 10. Corporate valuation model The corporate valuation model, the price-to-earnings (P/E) multiple approach, and the economic value-added (EVA) approach are some examples of valuation techniques. The corporate valuation model is similar to the dividend-based valuation that youve done in previous problems, but it focuses on a firms free instead of its dividends. Some firms dont pay dividends, or their dividends are difficult to forecast. For reason, some analysts use the corporate valuation model. cash flows (FCFs) that Tropetech Inc. has an expected net operating profit after taxes, EBIT(1 T), of $10,200 million in the coming year. In addition, the firm is expected to have net capital expenditures of $1,530 million, and net operating working capital (NOWC) is expected to increase by $40 million. How much free cash flow (FCF) is Tropetech Inc. expected to generate over the next year? O $128,693 million $8,710 million $8,630 million ?$11,690 million Tropetech Inc.s FCFs are expected to grow at a constant rate of 5.70% per year in the future. The market value of Tropetech Inc.s outstanding debt is $34,066 million, and preferred stocks value is $18,925 milion. Tropetech Inc. has 600 million shares of common stock outstanding, and its weighted average cost of capital (WACC) equals 17.10%. Using the preceding information and the FCE you calculated in the previous question, calculate the appropriate values in this table Term Value (Millions) Total firm value Value of common equity Intrinsic value per share
Answer #1

1)correct option is "C" -8630 million

Free cash flow = EBIT [1-TAX] -capital expenditure -increase in working capital

          = 10200-1530-40

          = 8630

2)Total Firm value = FCF /(WACC-g)

            = 8630 /(.1710 -.057)

           = 8630 / .114

           = $ 75701.75

value of equity =value of firm -value of debt -value of preferred stock

         = 75701.75- 34066-18925

           = 22710.75

Intrisic value per share = 22710.75 / 600 = $ 37.85 per share